Assume that Ocean King Products sells three varieties of canned seafood with the following prices and costs.

Selling Price
per Case
Variable Cost
per Case
Fixed Cost
per Month
Variety 1 $ 10 $ 8
Variety 2 8 6
Variety 3 13 9
Entire firm $ 46,800

The sales mix (in cases) is 40 percent Variety 1, 35 percent Variety 2, and 25 percent Variety 3.

Required:

a. At what sales revenue per month does the company break even?

b. Suppose the company is subject to a 35 percent tax rate on income. At what sales revenue per month will the company earn $43,485 after taxes assuming the same sales mix?
Answer #1

Ans:

Ocean King Products:

Products Selling price per case $ Variable cost per case $ Contribution margin per case $
Variety 1 10 8 2
Variety 2 8 6 2
Variety 3 13 9 4

Sales mix: variety 1 = 40%, variety 2 = 35%, variety 3 = 25%

Let total sales be x

Contribution margin:

Variety 1 = 2 * x * 40% = $ 0.8x

Variety 2 = 2 * x * 35% = $ 0.7x

Variety 3 = 4 * x * 25% = $ 1x

Total contribution margin = $ 2.5x

Break even sales units = Fixed cost / contribution margin

= $46,800/2.5x

x = 18,720

Break even sales revenue: (Units sold * Sales mix * Selling price)

Variety 1 = 18,720 * 40% * $10 = $74,880

Variety 2 = 18,720 * 35% * $8 = $52,416

Variety 3 = 18,720 * 25% * $13 = $60,840

a. Total break even sales revenue = $188,136

b. Profit before taxes = $43,485 / 65%

Profit before taxes = $66,900

Required sales = (Required profit + Fixed cost) / Contribution margin

= ($66,900 + $46,800) / 2.5x

x = $113,700/2.5

x = 45,480 units to be sold for profit of $43,485 after taxes

Sales revenue before taxes:

Variety 1 = 45,480 * $10 * 40% = $181,920

Variety 2 = 45,480 * $8 * 35% = $127,344

Variety 3 = 45,480 * $13 * 25% = $147,810

Total sales revenue= $457,074 to earn profit of $43,485 after taxes

b. Total sales revenue =$457,074

a. Break even sales revenue $188,136
b. Sales revenue to earn $43,485 after taxes $457,074

If any doubts or queries please comment and clarify I'll explain ASAP.